# Gompertz Linear Units: Leveraging Asymmetry for Enhanced Learning Dynamics

### v1

@@ Abstract
Activation functions are fundamental elements of deep learning architectures as they significantly influence training dynamics. ReLU, while widely used, is prone to the dying neuron problem, which has been mitigated by variants such as LeakyReLU, PReLU, and ELU that better handle negative neuron outputs. Recently, self-gated activations like GELU and Swish have emerged as state-of-the-art alternatives, leveraging their smoothness to ensure stable gradient flow and prevent neuron inactivity. In this work, we introduce the Gompertz Linear Unit (GoLU), a novel self-gated activation function defined as $\mathrm{GoLU}(x)=x\,\mathrm{Gompertz}(x)$, where $\mathrm{Gompertz}(x)=e^{-e^{-x}}$. The GoLU activation leverages the asymmetry in the Gompertz function to reduce variance in the latent space more effectively compared to GELU and Swish, while preserving robust gradient flow. Extensive experiments across diverse tasks, including Image Classification, Language Modeling, Semantic Segmentation, Object Detection, Instance Segmentation, and Diffusion, highlight GoLU’s superior performance relative to state-of-the-art activation functions, establishing GoLU as a robust alternative to existing activation functions.
@@ 2.1 Definition and Properties
In this section, we introduce the GoLU activation function and discuss its properties. GoLU is defined through Equations 1 and 2 and visualized in Figure 1 (Left).
@@
The gate function $\mathrm{Gompertz(x)}$ refers to the Gompertz function introduced in (Gompertz, 1825) and is plotted in Figure 1 (Right).
Figure 1: GoLU activation (Left) and its Gompertz gate (Right). The Gompertz gate depicts a slight rightward shift.
@@
From Equations 1, 2 and Figure 1, we understand that, contrary to ReLU and its variants which are monotonic and non-smooth at zero, GoLU is a smooth and non-monotonic self-gated activation, similar to Swish and GELU. In fact the formulation of GoLU using exponentials makes it infinitely differentiable. However, in contrast to Sigmoid and the Gaussian CDF (i.e. the gate functions of Swish and GELU), the Gompertz function is asymmetric, as it does not mirror evenly around a central point. This asymmetry, which has a bias towards the right, arises from the inherent asymmetry of the Gumbel distribution, which favors positive input values. In fact, the right-leaning asymmetry of the Gumbel distribution leads to smaller gate values across the entire input range, inducing a compression effect on the output distribution. This behavior extends to GoLU, yielding output values closer to zero, both for positive and negative inputs, when compared to other gated activation functions, effectively reducing the magnitude of the activation output. We note that, while Mish also exhibits an asymmetric distribution, it is skewed to the left, producing the opposite effect relative to GoLU.
These properties are more clearly illustrated in Figure 2, which provides a direct comparison between different activations (Left), as well as the gate functions of various gated activations (Middle) and their corresponding distributions (Right).
Additionally, from a more localized perspective, the Gompertz gate exhibits a reduced value in particular at the origin. This leads to a decreased steepness of GoLU near this point, as indicated by $\mathrm{GoLU}^{\prime}(0)=\mathrm{Gompertz}(0)$ from Equation 1. This property of reduced slope magnitude is not confined to the origin but extends to a neighborhood around it and spans a substantial portion of the negative input domain. Additional details are provided in Appendix A.
@@
Refer to captionRefer to caption Refer to caption
Figure 2: Comparison of different activation functions (Left) gate functions for gated activations (Middle) and the corresponding distributions (Right).
@@ 2.2 Effects on Training Dynamics
As illustrated in Figure 2 (Left), GoLU exhibits a profile that remains closest to the x-axis across the entire input range. Moreover, its slope, particularly near the origin and over a substantial portion of the negative input domain, is smaller in magnitude compared to other gated activations, as pointed out in Section 2.1. These characteristics suggest a reduced sensitivity of the activation output to variations in the input. This effectively reduces variance in the latent representations, and promotes smoother activation outputs, enhancing the model’s ability to differentiate between strong and weak features.
@@
Reduced activation variance results in less noisy and more consistent gradients. This typically means that the loss function changes more smoothly with respect to model parameters. As a result, the optimizer is more likely to converge to flatter regions of the loss landscape with smaller curvature. This is expected to result in better robustness to small perturbations of the model parameters. We explore this by adding two different Standard Normal noise terms, scaled independently by $\alpha,\beta$, to the weights of ResNet-20 trained on CIFAR-10. We compute the test loss across a grid of scaling factors $\alpha,\beta$ for the two terms, while keeping the noises constant (refer to Appendix B for more details). ResNet-20 with GoLU shows relatively smoother, less-spiked loss landscapes compared to other activations (Figure 5) which implies better generalization and noise robustness with GoLU. In contrast, ReLU’s non-smooth nature produces a highly-spiked landscape.
Figure 5: The loss landscape on the test set of ResNet-20 trained on CIFAR-10 with ReLU, GELU, Swish and GoLU after adding random, scaled perturbations to the learned weights (refer to Appendix B for more details).
@@ 3.1 Overview of Experiments
In Appendix D we further present a Critical Difference analysis to systematically compare the overall performance of activation functions. Finally, in Appendix G, we explore the application of GoLU to the task of learning curve extrapolation.
@@ 3.2 Image Classification
Table 2 presents the top-1 test accuracies with standard errors for ResNets 18, 34 and 50, WideResNet-50-2, DenseNet-121, EfficientNet-B0, ViT-B/32, ViT-B/16 and TinyViT (Wu et al., 2022). The training settings, detailed in Appendix F.1, are adopted from Torchvision (TorchVision, 2016) for all experiments except EfficientNet-B0 which is taken from the timm library (Wightman, 2019) and TinyViT which is taken from (Wu et al., 2022).
As highlighted, GoLU consistently outperforms the standard activation functions across all architectures, with the exception of EfficientNet-B0, where the performance difference is minimal. Notice that EfficientNet-B0 is an exception because its nonlinearity arises not only from activation functions (which are replaced) but also from a squeeze-and-excitation block, which remains unchanged in our experiments. For ResNet-50 and ViT-B/32, test loss and test accuracy curves are shown in Figures 7 and 8, respectively, where GoLU consistently delivers lower test loss and higher top-1 accuracy over the epochs. GELU is generally the second-best performer, while ELU performs worst across most architectures.
@@ 3.3 Language Modeling
We train babyGPT on TS and GPT2-S (124M) on OWT, both sourced from the nanoGPT repository (Karpathy, 2023). As shown in Table 4, GoLU demonstrates superior performance, achieving lower perplexity and higher token accuracy on both babyGPT and GPT2-S. GoLU’s superiority is also evident in the test loss curves in Figures 9 and 10. The general trend of GELU being the second-best activation function holds in language modeling as well.
@@
Appendix F.3 outlines the architectural details and provides additional information on the datasets and training settings.
@@ 3.4 Semantic Segmentation
For Semantic Segmentation, we train DeepLabV3 on the MS-COCO dataset with PASCAL-VOC labels, from the Torchvision benchmark (see Appendix F.4). We employ our ResNet-50 backbone, pre-trained on ImageNet-1k.
Table 5 presents the test loss and test mIoU using the original learning rate of 0.02. GoLU achieves the lowest test loss, whereas ReLU attains the highest mIoU, with GoLU ranking second. However, the difference in mIoU between ReLU and GoLU is statistically insignificant.
@@
We conduct a small ablation study on the learning rate and find that lr=0.02 is suboptimal for training the model. Instead, lr=0.01 yields the best performance across all activation functions (see heatmap 19 in Appendix C for full results). Table 5 also reports the results with lr=0.01, where GoLU consistently outperforms other activation functions in terms of mIoU. Additionally, the inference loss and test mIoU curves over epochs, shown in Figures 11 and 12, further emphasize GoLU’s strong performance in semantic segmentation.
@@ 3.5 Object Detection
For Object Detection, we train Faster R-CNN-FPN and RetinaNet-FPN on the MS-COCO dataset. As shown in Table 6 and Figure 13, GoLU outperforms all activation functions on object detection as well, with higher Box mAP (AP @ IoU=0.50:0.95, area=all, maxDets=100) across both Faster R-CNN-FPN and RetinaNet-FPN architectures, while GELU ranks second. Appendix F.5 outlines experimental details.
@@
Activation Faster R-CNN Box mAP RetinaNet Box mAP
@@ 3.6 Instance Segmentation
For Instance Segmentation, we train Mask R-CNN-FPN with a ResNet-50 backbone from the Torchvision benchmark on the MS-COCO dataset (see Appendix F.6 for training settings). As shown in Table 7, GELU achieves the best performance in this setting, with GoLU ranking second in Box mAP and third in Mask mAP (both implying AP @ IoU=0.50:0.95, area=all, maxDets=100). However, Figure 14, which depicts test Box mAP and Mask mAP over epochs, reveals that GoLU generally outperforms GELU and ReLU throughout the training process. Based on these observations, we suggest that, similar to the Semantic Segmentation task, the learning rate of 0.02 may be suboptimal for this specific architecture-dataset combination. Adjusting the learning rate could potentially enhance GoLU’s performance relative to baseline activations.
@@
Activation Box mAP Mask mAP
ReLU 38.33±0.001 34.19±0.001
LeakyReLU 38.31±0.002 34.19±0.001
ELU 36.41±0.001 32.81±0.001
GELU 39.00±0.001 34.73±0.000
Swish 38.19±0.002 33.99±0.001
Mish 38.76±0.000 34.70±0.000
GoLU 38.96±0.001 34.54±0.001
@@ 3.7 Denoising Diffusion Probabilistic Models
We train a Denoising Diffusion Probabilistic Model on the CelebA dataset (see Appendix F.7). As shown in Table 8, for the default lr=0.0003, gated activations perform comparably to the baseline activation, Swish, which achieves the best performance, with GoLU ranking a close second. Figure 15 (Left) further illustrates the test loss over epochs. Similar to our findings in semantic segmentation, we conduct a learning rate ablation study. Results, summarized in heatmap 19 in Appendix C, indicate that increasing the lr from the default value of 0.0003 to 0.0004, 0.0005 and 0.001 progressively improves performance across all activations. Notably, for lr values of 0.0004, 0.0005 and 0.001, GoLU achieves the lowest final test loss. Results for the optimum lr=0.001 are highlighted in the right column of Table 8 and Figure 15 (Right). These findings are in line with the trend observed in semantic segmentation, where GoLU outperforms baseline activations under optimal lr configurations.
@@
Activation Test Loss Test Loss
@@ 4 Training and Inference Speed
Existing activation functions in PyTorch leverage CUDA kernels in Eager mode to achieve optimal speedup. To ensure a fair comparison of training and inference speeds, we developed a CUDA-optimized kernel for GoLU, which was used for all training experiments described in the previous sections. Table 9 in Appendix E presents the relative training and inference speeds of GoLU compared to the default activation function across various tasks.
@@ Acknowledgements
This research was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828. The authors gratefully acknowledge the computing time made available to them on the high-performance computer NHR@KIT Compute Cluster at the NHR Center NHR@KIT. These Centers are jointly supported by the Federal Ministry of Education and Research and the state governments participating in the NHR (www.nhr-verein.de/unsere-partner). We acknowledge funding by the European Union (via ERC Consolidator Grant DeepLearning 2.0, grant no. 101045765). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Center (JSC) and SuperMUC-NG at Leibniz Supercomputing Centre (www.lrz.de). The authors acknowledge support by the state of Baden-Württemberg through bwHPC. Frank Hutter acknowledges the financial support of the Hector Foundation. We also thank Jörg Franke, Edward Bergman, Arbër Zela, André Biedenkapp and Lennart Purucker for their constructive feedback throughout the development of this work.

### v2

@@ Abstract
Activation functions are fundamental elements of deep learning architectures as they significantly influence training dynamics. ReLU, while widely used, is prone to the dying neuron problem, which has been mitigated by variants such as LeakyReLU, PReLU, and ELU that better handle negative neuron outputs. Recently, self-gated activations like GELU and Swish have emerged as state-of-the-art alternatives, leveraging their smoothness to ensure stable gradient flow and prevent neuron inactivity. In this work, we introduce the Gompertz Linear Unit (GoLU), a novel self-gated activation function defined as $\mathrm{GoLU}(x)=x\,\mathrm{Gompertz}(x)$, where $\mathrm{Gompertz}(x)=e^{-e^{-x}}$. The GoLU activation leverages the right-skewed asymmetry in the Gompertz function to reduce variance in the latent space more effectively compared to GELU and Swish, while preserving robust gradient flow. Extensive experiments across diverse tasks, including Image Classification, Language Modeling, Semantic Segmentation, Object Detection, Instance Segmentation, and Diffusion, highlight GoLU’s superior performance relative to state-of-the-art activation functions, establishing GoLU as a robust alternative to existing activation functions.
@@ 2.1 Definition and Properties
In this section, we introduce the GoLU activation function and discuss its properties. GoLU is defined through Equations 1 and 2 and visualized in Figure 1 (Left) as the red curve, alongside other activation functions for comparison.
@@
The gate function $\mathrm{Gompertz(x)}$ refers to the Gompertz function introduced in (Gompertz, 1825) and is plotted in red in Figure 1 (Right), alongside the gate functions of other gated activations.
Figure 1: Activation functions (Left) and their corresponding gate functions (Right). GoLU and its gate, the Gompertz function, are highlighted in red. Note the slight rightward shift of the Gompertz gate.
@@
From Equations 1, 2 and Figure 1, we understand that, contrary to ReLU and its variants which are monotonic and non-smooth at zero, GoLU is a smooth and non-monotonic self-gated activation, similar to Swish and GELU. In fact the formulation of GoLU using exponentials makes it infinitely differentiable. However, in contrast to Sigmoid and the Gaussian CDF (i.e. the gate functions of Swish and GELU), the Gompertz function is asymmetric, as it does not mirror evenly around a central point [^1] Formally, we refer to a scalar function $f(x)$ as symmetric if there exists a point $x^{*}$ such that for any input $x$ we have $f(x^{*}+x)-f(x^{*})=f(x^{*})-f(x^{*}-x)$..
This asymmetry, which has a bias towards the right, arises from the inherent asymmetry of the Gumbel distribution, which favors positive input values, as illustrated in Figure 2 (Left). In fact, the right-leaning asymmetry of the Gumbel distribution leads to smaller gate values across the entire input range, inducing a compression effect on the output distribution. This behavior extends to GoLU, yielding output values closer to zero, both for positive and negative inputs, when compared to other gated activation functions, effectively reducing the magnitude of the activation output. We note that, while Mish also exhibits an asymmetric distribution, it is skewed to the left, producing the opposite effect relative to GoLU [^2] See Appendix B for an interesting case of flipped Mish distribution with right-leaning asymmetry..
From a more localized perspective, the Gompertz gate exhibits a reduced value in particular at the origin. This leads to a decreased steepness of GoLU near this point, as indicated by $\mathrm{GoLU}^{\prime}(0)=\mathrm{Gompertz}(0)$ from Equation 1. This property of reduced slope magnitude is not confined to the origin but extends to a neighborhood around it and spans a substantial portion of the negative input domain, as shown in Figure 2 (Right). Additional details are provided in Appendix A.
@@
Figure 2: Comparison of the distributions underlying the gate functions (Left) and the gradients (Right) of various gated activations. The Gumbel distribution exhibits a slight rightward skew.
@@ 2.2 Effects on Training Dynamics
As illustrated in Figure 1 (Left), GoLU exhibits a profile that remains closest to the x-axis across the entire input range. Moreover, its slope, particularly near the origin and over a substantial portion of the negative input domain, is smaller in magnitude compared to other gated activations, as pointed out in Section 2.1. These characteristics suggest a reduced sensitivity of the activation output to variations in the input. In fact, for a scalar activation function $f$ the variance of its output can be shown to be approximately proportional to the square of its slope
$\text{Var}[f(x)]\approx f^{\prime}(\mu)^{2}\sigma^{2}$ (4)
where $\mu$ and $\sigma^{2}$ denote the mean and variance of the input, respectively (see Appendix A for the derivation). This analytic relation explains more directly how the smaller slope of GoLU contributes to a lower variance in its output. As a result, GoLU effectively reduces variance in the latent representations, and promotes smoother activation outputs, enhancing the model’s ability to distinguish between strong and weak features.
@@
Reduced activation variance results in less noisy and more consistent gradients. This typically means that the loss function changes more smoothly with respect to model parameters. As a result, the optimizer is more likely to converge to flatter regions of the loss landscape with smaller curvature. This is expected to result in better robustness to small perturbations of the model parameters. We explore this by adding two different Standard Normal noise terms, scaled independently by $\alpha,\beta$, to the weights of ResNet-20 trained on CIFAR-10. We compute the test loss across a grid of scaling factors $\alpha,\beta$ for the two terms, while keeping the noises constant (refer to Appendix C for more details). ResNet-20 with GoLU shows relatively smoother, less-spiked loss landscapes compared to other activations (Figure 5) which implies better generalization and noise robustness with GoLU. Additionally, GoLU exhibits a lower average loss, as shown in Figure 5.
Figure 5: The loss landscape on the test set of ResNet-20 trained on CIFAR-10 with ReLU, GELU, Swish and GoLU after adding random, scaled perturbations to the learned weights (refer to Appendix C for more details).
Figure 9 in Appendix C presents a comparison of loss value distributions across the loss landscape, indicating that GoLU yields lower loss variance compared to other activation functions.
@@ 3.1 Overview of Experiments
In Appendix E we further present a Critical Difference analysis to systematically compare the overall performance of activation functions. In Appendix I, we assess the performance of GoLU on a machine translation task, using the WMT14 English–German benchmark. Finally, in Appendix J, we explore the application of GoLU to the task of learning curve extrapolation.
@@ 3.2 Image Classification
Table 2 presents the top-1 test accuracies with standard errors for ResNets 18, 34 and 50, WideResNet-50-2, DenseNet-121, EfficientNet-B0, ViT-B/32, ViT-B/16 and TinyViT (Wu et al., 2022). The training settings, detailed in Appendix G.1, are adopted from Torchvision (TorchVision, 2016) for all experiments except EfficientNet-B0 which is taken from the timm library (Wightman, 2019) and TinyViT which is taken from (Wu et al., 2022).
As highlighted, GoLU consistently outperforms the standard activation functions across all architectures, with the exception of EfficientNet-B0, where the performance difference is minimal. Notice that EfficientNet-B0 is an exception because its nonlinearity arises not only from activation functions (which are replaced) but also from a squeeze-and-excitation block, which remains unchanged in our experiments. For ResNet-50 and ViT-B/32, test loss and test accuracy curves are shown in Figures 17 and 18, respectively, where GoLU consistently delivers lower test loss and higher top-1 accuracy over the epochs. GELU is generally the second-best performer, while ELU performs worst across most architectures.
@@ 3.3 Language Modeling
We train babyGPT on TS and GPT2-S (124M) on OWT, both sourced from the nanoGPT repository (Karpathy, 2023). As shown in Table 4, GoLU demonstrates superior performance, achieving lower perplexity and higher token accuracy on both babyGPT and GPT2-S. GoLU’s superiority is also evident in the test loss curves in Figures 19 and 20. The general trend of GELU being the second-best activation function holds in language modeling as well.
@@
Appendix G.3 outlines the architectural details and provides additional information on the datasets and training settings.
@@ 3.4 Semantic Segmentation
For Semantic Segmentation, we train DeepLabV3 on the MS-COCO dataset with PASCAL-VOC labels, from the Torchvision benchmark (see Appendix G.4). We employ our ResNet-50 backbone, pre-trained on ImageNet-1k.
Table 5 (Left) presents the test loss and test mIoU using the original learning rate of 0.02. GoLU achieves the lowest test loss, whereas ReLU attains the highest mIoU, with GoLU ranking second. However, the difference in mIoU between ReLU and GoLU is statistically insignificant.
@@
We conduct a small ablation study on the learning rate and find that lr=0.02 is suboptimal for training the model. Instead, lr=0.01 yields the best performance across all activation functions (see heatmap 11 in Appendix D for full results). Table 5 (Right) reports the results with lr=0.01, where GoLU consistently outperforms other activation functions in terms of mIoU. Additionally, the inference loss and test mIoU curves over epochs, shown in Figures 21 and 22, further emphasize GoLU’s strong performance in semantic segmentation.
@@ 3.5 Object Detection
For Object Detection, we train Faster R-CNN-FPN and RetinaNet-FPN on the MS-COCO dataset. As shown in Table 6 and Figure 23, GoLU outperforms all activation functions on object detection as well, with higher Box mAP (AP @ IoU=0.50:0.95, area=all, maxDets=100) across both Faster R-CNN-FPN and RetinaNet-FPN architectures, while GELU ranks second. Appendix G.5 outlines experimental details.
@@
Activation Box mAP
Faster R-CNN RetinaNet
@@ 3.6 Instance Segmentation
For Instance Segmentation, we train Mask R-CNN-FPN with a ResNet-50 backbone from the Torchvision benchmark on the MS-COCO dataset (see Appendix G.6 for training settings). As shown in Table 7 (Left), GELU achieves the best performance in this setting (with the default lr=0.02), with GoLU ranking second in Box mAP and third in Mask mAP (both implying AP @ IoU=0.50:0.95, area=all, maxDets=100). However, Figure 24, which depicts test Box mAP and Mask mAP over epochs, reveals that GoLU generally outperforms GELU and ReLU throughout the training process. We perform a learning rate ablation study and observe that, similar to the Semantic Segmentation task, a learning rate of 0.02 is suboptimal for this specific architecture–dataset combination. In contrast, increasing the learning rate to 0.03 leads to improved performance across all activation functions (see heatmaps 13 and 13). Surprisingly, at the optimal learning rate of 0.03, GoLU outperforms all baseline activations, as shown in Table 7 (right).
@@
Activation LR=0.02 LR=0.03
Box mAP Mask mAP Box mAP Mask mAP
ReLU 38.33±0.133 34.19±0.129 38.25±0.144 34.28±0.116
LeakyReLU 38.31±0.230 34.19±0.160 38.42±0.105 34.26±0.065
ELU 36.41±0.087 32.81±0.082 35.85±0.069 32.38±0.043
GELU 39.00±0.073 34.73±0.024 39.12±0.028 34.93±0.061
Swish 38.19±0.224 33.99±0.168 38.10±0.187 33.95±0.112
Mish 38.76±0.036 34.70±0.026 38.56±0.138 34.64±0.078
GoLU 38.96±0.101 34.54±0.083 39.36±0.192 34.97±0.146
@@ 3.7 Denoising Diffusion Probabilistic Models
We train a Denoising Diffusion Probabilistic Model on the CelebA dataset (see Appendix G.7). As shown in Table 8, for the default lr=0.0003, gated activations perform comparably to the baseline activation, Swish, which achieves the best performance, with GoLU ranking a close second. Figure 25 (Left) further illustrates the test loss over epochs. Similar to our findings in semantic segmentation and instance segmentation, we conduct a learning rate ablation study. Results, summarized in heatmap 11 in Appendix D, indicate that increasing the lr from the default value of 0.0003 to 0.0004, 0.0005 and 0.001 progressively improves performance across all activations. Notably, for lr values of 0.0004, 0.0005 and 0.001, GoLU achieves the lowest final test loss. Results for the optimum lr=0.001 are highlighted in the right column of Table 8 and Figure 25 (Right). These findings are in line with the trend observed in semantic segmentation and instance segmentation, where GoLU outperforms baseline activations under optimal lr configurations.
@@
Activation Test Loss
@@ 4 Training and Inference Speed
Existing activation functions in PyTorch leverage CUDA kernels in Eager mode to achieve optimal speedup. To ensure a fair comparison of training and inference speeds, we developed a CUDA-optimized kernel for GoLU, which was used for all training experiments described in the previous sections. Table 9 in Appendix F presents the relative training and inference speeds of GoLU compared to the default activation function across various tasks.
@@ Acknowledgements
This research was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 539134284, through EFRE (FEIH_2698644) and the state of Baden-Württemberg.
[Uncaptioned image][Uncaptioned image]
The authors gratefully acknowledge the computing time made available to them on the high-performance computer NHR@KIT Compute Cluster at the NHR Center NHR@KIT. These Centers are jointly supported by the Federal Ministry of Education and Research and the state governments participating in the NHR (www.nhr-verein.de/unsere-partner). We acknowledge funding by the European Union (via ERC Consolidator Grant DeepLearning 2.0, grant no. 101045765). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Center (JSC) and SuperMUC-NG at Leibniz Supercomputing Centre (www.lrz.de). The authors acknowledge support by the state of Baden-Württemberg through bwHPC. We acknowledge the EuroHPC Joint Undertaking for awarding this project access to the EuroHPC supercomputer LEONARDO, hosted by CINECA (Italy) and the LEONARDO consortium through an EuroHPC Extreme Access grant EHPC-EXT-2023E02-068. Frank Hutter acknowledges the financial support of the Hector Foundation. We also thank Jörg Franke, Edward Bergman, Arbër Zela, André Biedenkapp and Lennart Purucker for their constructive feedback throughout the development of this work.